OpenStudy (anonymous):

Question below:

OpenStudy (anonymous):

Let \[\left\{ u _{1}, u_{2} \right\} \space and \space \left\{ v_{1},v_{2} \right\}\] be ordered basis for R^2 where \[u_{1} = \left(\begin{matrix}1 \\ 1\end{matrix}\right)\] , \[u_{2} = \left(\begin{matrix}-1 \\ 1\end{matrix}\right)\] and \[v_{1} = \left(\begin{matrix}2 \\ 1\end{matrix}\right)\] , \[v_{2} = \left(\begin{matrix}1 \\ 0\end{matrix}\right)\] Let L be the linear transformation defined by \[L(x) = (-x_{1},x_{2})\] and let B be the matrix representing L wrt [u1,u2} a) Find the transition matrix from change of basis from U to V b) Find the matrix A representing L wrt V by computing \[S^{-1}AS\]

OpenStudy (anonymous):

I've got B to be \[B = \left[\begin{matrix}-1 & 1 \\ 1 & 1\end{matrix}\right]\]

OpenStudy (anonymous):

@amistre64 @myininaya @hartnn

OpenStudy (anonymous):

@Directrix @Zarkon

OpenStudy (anonymous):

Should be \[S^{-1}BS\]

OpenStudy (anonymous):

@Hero ?

OpenStudy (anonymous):

@terenzreignz ??

terenzreignz (terenzreignz):

First, express \(v_1\) and \(v_2\) as linear combinations of \(u_1\) and \(u_2\)

terenzreignz (terenzreignz):

Wait, topsy turvy, express u1 and u2 at linear combinations of v1 and v2 D:

OpenStudy (anonymous):

u1 = v1 - v2 u2 = v1 - 3v2

OpenStudy (anonymous):

So that makes s \[S = \left[\begin{matrix}1 & -1 \\ 1 & -3\end{matrix}\right]\]

terenzreignz (terenzreignz):

No, you arrange them vertically XD \[\Large S^\color{red}T = \left[\begin{matrix}1 & -1 \\ 1 & -3\end{matrix}\right]\]

OpenStudy (anonymous):

Oh yeah, i knew that xD

OpenStudy (anonymous):

\[S = \left[\begin{matrix}1 & 1 \\ -1 & -3\end{matrix}\right]\]

OpenStudy (anonymous):

So \[S^{-1} = -\frac{1}{2}\left[\begin{matrix}-3 & -1 \\ 1 & 1\end{matrix}\right]\]

terenzreignz (terenzreignz):

Okay, sure... (I'll take your word for that)

OpenStudy (anonymous):

Yeah its right, i checked :P now i just gotta multiply all this together D:

OpenStudy (anonymous):

I got it from here, thank you again for your help! :D

terenzreignz (terenzreignz):

Awesome. No problem :D